# CoDiff: Conditional Diffusion Model for Collaborative 3D Object Detection

### v2

@@ V-C Quantitative evaluation
TABLE II: Selection of the compression rate $\tau$ in CoDiff.
@@
TABLE III: Ablation study on V2XSet val set.
@@
TABLE IV: Performance comparison of different sampling methods at varying steps in the proposed CoDiff.
@@
TABLE V: Inference time and Parameters.
@@
V-D Qualitative evaluation
Detection visualization. We show the 3D detection results in the Bird’s-eye-view (BEV) format on the V2XSet. Red and green boxes denote the detection results and the ground-truth, respectively. The degree of overlapping of these boxes reflects the performance of a method. Figure 4 depicts the detection results of V2X-ViT, CoAlign, RoCo and the proposed CoDiff at an intersection to validate the effectiveness of our method. We set the noise level and time delay of $0.4m/0.4^{\circ}$ and 100ms to produce high noises to make the perception task challenging. From the figure, it can be observed that V2X-ViT has many missed detections, while CoAlign [21] and RoCo [17] generate many predictions with relatively large offsets. In contrast, our CoDiff demonstrates strong performance in the presence of pose errors and time delays.
V-E Ablation Studies
To determine the optimal compression rate $\tau$ during the perceptual compression, we conduct ablation experiments on V2XSet dataset, as shown in Table II. We found that setting $\tau$ to 32 and compressing the feature dimensions to 8 achieved the optimal mAP. To validate the effectiveness of the conditional diffusion module, Table III shows the results of using or not using conditions, as well as how the conditions were applied in the model. Please note: (i) The diffusion conditions enhance the stability of the model; (ii) Using Concat for condition fusion is superior to directly applying element-wise addition. In testing, we used two samplers: DDPM [44] and DDIM [45]. To achieve optimal performance, we tested the effect of different samplers and sampling steps on the results. The experimental results are shown in Table IV. We selected DDPM with 8 sampling steps as the default solver in the experiments due to its simplicity and efficiency. We also measure the inference time and the parameters of our proposed CoDiff on a single NVIDIA 2080Ti GPU in Table V. At 1-step DDPM sampling, CoDiff can reach the inference time of 127.7 (ms) with 769.17 (MB) parameters.
@@ VI CONCLUSION
This paper proposes a novel robust collaborative perception framework CoDiff for 3D object detection. This framework leverages diffusion models to address information inaccuracies caused by pose errors and time delays between multiple agents, enhancing the precision and reliability of information exchange. Comprehensive experiments demonstrate that CoDiff achieves outstanding performance across all settings and exhibits exceptional robustness under extreme noise conditions. In future work, we aim to further explore the potential of CoDiff and expand its application to a broader range of perception tasks.
ACKNOWLEDGMENT
This work was supported by the National Natural Science Foundation of China Grant No. 12071478, No. 61972404; Public Computing Cloud and the Blockchain Lab, School of Information, Renmin University of China.

### v3

@@ V-C Quantitative evaluation
TABLE II: Component ablation study.
Modules AP@0.5 AP@0.7
PCM TCM 0.2/0.2 0.4/0.4 0.2/0.2 0.4/0.4
× × 73.8 72.0 58.8 57.9
× 74.8 73.3 59.4 58.4
× 75.0 72.1 59.8 57.4
75.4 73.6 60.0 59.0
V-D Qualitative evaluation
Detection visualization. We show the 3D detection results in the Bird’s-eye-view (BEV) format on the V2XSet. The degree of overlapping of these boxes reflects the performance of a method. Figure 6 depicts the detection results of V2X-ViT, CoAlign, RoCo and the proposed CoDiff at an intersection to validate the effectiveness of our method. We set the noise level and time delay of $0.4m/0.4^{\circ}$ and 100ms to produce high noises to make the perception task challenging. From the figure, it can be observed that V2X-ViT has many missed detections, while CoAlign [23] and RoCo [19] generate many predictions with relatively large offsets. In contrast, our CoDiff demonstrates strong performance in the presence of pose inconsistencies and time delays.
TABLE III: Selection of the compression rate $\tau$ in CoDiff.
@@ V-D Qualitative evaluation
TABLE IV: Selection of the threshold value $\tau_{2}$ in Eq. (3).
DAIR-V2X V2XSet
Threshold/Metric AP@0.7
Noise Level 0.2/0.2 0.4/0.4 0.6/0.6 0.8/0.8 0.2/0.2 0.4/0.4 0.6/0.6 0.8/0.8
$\tau_{2}$ = 1 58.8 57.6 56.9 56.4 77.1 72.5 70.4 65.7
$\tau_{2}$ = 2 58.9 57.9 57.2 56.9 77.4 77.3 71.0 68.1
$\tau_{2}$ = 3 59.4 58.4 58.2 57.8 77.1 73.1 70.2 68.9
$\tau_{2}$ = 4 57.9 57.6 57.6 57.5 75.9 72.0 69.7 68.3
V-E Ablation Studies
Contribution of major components in CoDiff. Now we investigate the effectiveness of individual components in CoDiff. We evaluate the impact of each component by progressively adding PCM and TCM. As Table II demonstrates, all the modules are beneficial to the performance gains, among which the pose adjustment module is particularly helpful in enhancing detection accuracy under high-noise conditions.
Selection of the optimal compression rate $\tau$. The selection of the optimal compression rate in CoDiff is crucial, as it directly impacts the quality of information shared between agents and the overall detection performance. To determine the optimal compression rate $\tau$ during the perceptual compression, we conduct ablation experiments on V2XSet dataset, as shown in Table III. We found that setting $\tau$ to 32 and compressing the feature dimensions to 8 achieved the optimal mAP.
TABLE V: Contribution of graph similarity.
Modules AP@0.5 AP@0.7
Matching distance edge 0.2/0.2 0.4/0.4 0.2/0.2 0.4/0.4
× × × 73.8 72.0 58.8 57.9
× 73.0 71.0 58.8 58.0
× 73.3 72.1 59.0 58.1
74.8 73.3 59.4 58.4
TABLE VI: Ablation study on V2XSet val set.
@@ V-E Ablation Studies
Selection of the threshold value $\tau_{2}$. To identify the optimal value for the threshold $\tau_{2}$ in Eq. (3) during graph initialization, we perform ablation studies on multiple datasets, as shown in Table IV. On the DAIR-V2X dataset, we observe that setting $\tau_{2}$ to 3 yields the best detection accuracy. When $\tau_{2}$ is increased beyond this point, performance degrades—primarily because vehicles in real-world environments tend to keep safe distances, and an overly small threshold would hinder effective initial matching. We also conduct the same analysis on a simulated dataset. The results show that when the noise level is below $0.8\,\text{m}/0.8^{\circ}$, the best detection performance occurs at $\tau_{2}=2\,\text{m}$. However, as the noise level rises to $0.8\,\text{m}/0.8^{\circ}$, the optimal threshold shifts to $\tau_{2}=3$. This change is attributed to the characteristics of the V2XSet dataset, where vehicles are more numerous and densely packed, resulting in generally shorter inter-vehicle distances compared to real-world data.
Contribution of graph similarity. In Table V, we apply the similarity metrics defined in Eq. (7) and Eq. (8) to the object matching process on the DAIR-V2X dataset. To assess the individual contribution of each similarity measure to the overall graph matching performance, we progressively introduce (i) edge-based similarity and (ii) distance-based similarity. The first row in Table V serves as the baseline, where no similarity-based matching is applied. Experimental findings demonstrate that incorporating either form of similarity leads to notable improvements in detection accuracy.
TABLE VII: Performance comparison of different sampling methods at varying steps in the proposed CoDiff.
@@
TABLE VIII: Inference time and Parameters.
@@
Effectiveness of the conditional diffusion module. Table VI shows the results of using or not using conditions, as well as how the conditions were applied in the model. Please note that: (i) The conditions enhance the stability of our porposed method; (ii) Using Concat for condition fusion is superior to directly applying element-wise addition. In testing, we used two samplers: DDPM [49] and DDIM [50]. To achieve optimal performance, we tested the effect of different samplers and sampling steps on the results. The experimental results are shown in Table VII. We selected DDPM with 8 sampling steps as the default solver in the experiments due to its simplicity and efficiency. We also measure the inference time and the parameters of our proposed CoDiff on a single NVIDIA 2080Ti GPU in Table VIII. At 1-step DDPM sampling, CoDiff can reach the inference time of 127.7 (ms) with 769.17 (MB) parameters. At 1-step DDIM sampling, CoDiff can reach the inference time of 125.8 (ms) with 776.91 (MB) parameters.
@@ VI CONCLUSION
This paper proposes a novel robust collaborative perception framework for 3D object detection, called CoDiff, which leverages conditional diffusion models to address the problem of feature misalignment. CoDiff primarily consists of two modules: PCM and TCM. The PCM establishes reliable correspondences between objects and performs pose calibration through graph optimization, while the TCM employs a conditional diffusion model to remove delay-induced noise, thereby generating highly refined multi-agent fused features. CoDiff does not require any ground truth pose and time supervision, making it highly practical. Comprehensive experiments demonstrate that our method achieves outstanding performance across all settings and exhibits exceptional robustness under extreme noise conditions. In future work, we aim to further explore the potential of CoDiff and expand its application to a broader range of perception tasks.
Acknowledgments
This work was supported by the National Natural Science Foundation of China Grant No. 61972404, 12071478, the Public Computing Cloud of Renmin University of China, and the Blockchain Lab of the School of Information, Renmin University of China.